# Simulations of Global Solar Convection with a Fully Compressible CHORUS++ Code

### v2

@@ 2.1 Model parameters, initial values, and simulation setup
Density at $r=R_{\rm bot}$, $\varrho_{\rm top}$ $[g\cdot cm^{-3}]$ $2.1\cdot 10^{-1}$ $2.1\cdot 10^{-1}$
Density at $r=R_{\rm top}$, $\varrho_{\rm top}$ $[g\cdot cm^{-3}]$ $2.1\cdot 10^{-4}$ $1.05\cdot 10^{-2}$
@@
kinematic viscosity $\nu$ $[cm^{2}s^{-1}]$ $6.0\cdot 10^{12}$ $6.0\cdot 10^{13}$
entropy diffusion coef. $\kappa_{s}$ $[cm^{2}s^{-1}]$ $6.0\cdot 10^{13}$ $6.0\cdot 10^{13}$
$U^{\prime}_{\rm rms}$ [$\cdots$-1] $1.51\cdot 10^{4}$ $4.04\cdot 10^{4}$
@@
Ekman num. $\nu/(\Omega_{0}D^{2})$ $10^{-14}$ $4.853\cdot 10^{-3}$ $2.447\cdot 10^{-3}$
Taylor num. $4\Omega_{0}^{2}D^{4}/\nu^{2}$ $10^{19\sim 27}$ $1.698\cdot 10^{5}$ $6.682\cdot 10^{5}$
Prandtl num. $\nu/\kappa_{s}$ $10^{-6\sim-4}$ $0.1$ $1$
Rossby num. $U^{\prime}_{\rm rms}/(2\Omega_{0}D)$ $10^{-1\sim 0}$ $1.239\cdot 10^{-1}$ $1.433\cdot 10^{-2}$
The initial values are determined in the same manner as described in Chen et al. (2023), where the thermal heat conduction rate ($\kappa_{r}$) is given as a function of the heliocentric distance and the density scale ($N_{p}$; the ratio of the mass density at the upper and lower boundary surfaces) must be given. We used the same value of the thermal heat conduction rate and a different value of the density scale (because of the difference in the radius of the top boundary surface, $R_{\rm top}$).
Figure 1: Radial profile of the simulated plasma quantities in the solar convective shell: (a) mass density $\varrho$ and (c) temperature ($T$) at $t=1\cdot 10^{7}s$. In panels (b) and (d), the deviations of the simulated variable from the initial values, $\varrho_{0}$ and $T_{0}$, are given as $(\varrho/\varrho_{0}-1)$ and $(T/T_{0}-1)$, respectively. The deviation is less than $10^{-3}$ overall, except the density deviation reaches about $2\cdot 10^{-3}$ near the top boundary. In the bottom row, the radial profile of the square of buoyancy frequency ($N^{2}$) of the initial value and the buoyancy frequency averaged horizontally (latitudinally and longitudinally) for the state at $t=10^{7}s$ are given in panels (e) and (f), respectively. In these two plots, cross (+) marks are placed where $N^{2}>0$.
@@
The initial state is nearly stable except the near-surface regions at $r\sim R_{\rm top}$ where the plasma structure is unstable to perturbations or fluctuations. Figure 1(e) shows the radial profile of the square of the buoyancy frequency at the initial value,
@@
Figure 2 demonstrates the radial component of the plasma velocity ($V_{r}$) simulated at $t=1\cdot 10^{7}s$. The convective flows are seen in the simulated convection zone, as well as the convective cell structures on the top boundary surface. In Figure 1(f), the square of the buoyancy frequency $N^{2}$ is nearly all negative throughout the simulation volume at the selected moment.
Panels (a) and (c) of Figure 1, respectively, show the radial profile of the mass density and temperature averaged over each shell layer at $t=1\cdot 10^{7}$s. Panels (b) and (d), respectively, show the ratios of the simulated plasma density and temperature to those of the initial values. Although the simulated system will be in the convective state, the differences in plasma density and temperature are found to be very small. A notable difference is found in the density ratio near the top boundary surface as shown in panel (b); still, the deviation is 0. 25% in the largest.
3 Results
The kinetic energy density ($\varrho V^{2}/2$) is a good proxy to represent the entire simulated convective system. In the benchmark case (Chen et al., 2023), the kinetic energy density of the simulated stellar convection zone exhibits exponential growth and a brief overshoot/oversaturation, then reaches a plateau profile that corresponds to the statistically stable saturated convection state.
@@ 3 Results
Figure 3 shows the temporal profile of the kinetic energy density. With the actual solar parameters, the temporal profile of the kinetic energy density is different from that of the benchmark tests. In the logarithmic scale plot, the exponential growth of the energy density is found as a (near) linear line, roughly at $10\leq t\leq 20$ days. However, the overshooting and the following plateau profile (like those in Figure 7 of Chen et al., 2023) are not clearly seen. The current simulation run has been done only for $t\leq 1.5\cdot 10^{7}~{}{\rm s}\sim 170~{}{\rm d}$, but it is unclear whether or not the kinetic density will become more constant in an extended simulation run. Because our visual inspections did not find distinct differences in the figures for $t=1\cdot 10^{7}s$ and $t=1.5\cdot 10^{7}s$, we chose the simulated data at $t=10^{7}{\rm s}\sim 115{\rm d}$ as a moderately mature state of the convection properties.
3.1 Large-scale Convection/Granulation Cells
@@
Figure 4: The plasma quantities of the simulated solar convection zone at $t=10^{7}$ s in the Mollwide projection mapping. (a) The plasma temperature subtracted by the average ($7.4132\cdot 10^{4}$ K), (b) the radial component of the plasma flow ($V_{r}$), (c) the latitudinal component of the plasma flow ($V_{\theta}$), at $r=0.982\,R_{\sun}$. (e) the longitudinal component of the plasma flow ($V_{\phi}$). In the right column, panels (d) and (f) respectively show the radial component of the plasma flows at the middle depth of the simulated convection zone (at $r=0.846\,R_{\sun}$) and near the bottom boundary surface (at $r=0.709\,R_{\sun}$) in the same format as in panel (b). The colors are truncated at the largest absolute value so that the red (for positive values) and blue (for negative values) are assigned evenly across the zero values.
@@
As shown in Figure 1(b,d), the deviations of the simulated plasma variables from the initial values are relatively small in the radial direction. The temperature, averaged over the selected sphere, was $7.4132\cdot 10^{4}{\rm K}^{\circ}$, approximately 8 degrees higher than the initial value, and the largest deviation from the averaged state is about 6.7 degrees. The lower temperature structures are found at the boundaries of the convection cells that can be identified as negative velocity $V_{r}$ in Figure 4(b). The temperature and plasma motions near the surface are found to be rather uneven within convection cell structures. For example, the region of relatively high temperature (panel a) and large prograde plasma motion ($V_{\phi}$ in panel e) are found in the western (right) part of the convection cells near the equator (panel b). The latitudinal motion ($V_{\theta}$ in panel c) exhibits a weak but similar tendency to $V_{\phi}$; relatively large poleward flows are found near the boundary of each convective cell close to the poles in each hemisphere. In the right column in Figure 4(b,d,f), the distribution of $V_{r}$ at different depths is presented. The size and shape of the convection cells differ substantially at different depths, while several strong downward flows are found at the same locations over a wide range of depths.
3.2 Polar regions
@@
Figure 5: (a) the curl of the horizontal flows, (b) — (d) The $x$, $y$, and $z$ components of the simulated plasma flows around the north pole.
@@
Figure 5 offers the north pole view of the simulated plasma flows near the top boundary surface (at $r=0.982\,R_{\sun}$). Because the CHORUS++ code uses an unstructured grid system, issues due to the geometric singularity like that in the case of the spherical coordinate presentation are not present, and the plasma flows crossing the solar rotation axis (or $z$-axis) are treated naturally. The snapshot shown (at $t=1\cdot 10^{7}$ s) does not show clear polar vortex cell structures (e.g., discussed by Nagashima et al., 2011; Dikpati et al., 2024). By tuning our future simulation settings, we plan to investigate in more detail the formation of plasma-flow structures in the high-latitude and polar regions that are believed to play a major role in solar-cycle activities.
3.3 Meridional Circulation and Rotation
@@
Figure 6: Meridional cross-section view of the simulated plasma flows at $t=10^{7}s$. From left, the radial component of the plasma flow ($V_{r}$), the latitudinal component ($V_{\theta}$), and the rotation rate relative to the reference frame ($\omega:=V_{\phi}/(r\sin\theta$) are shown. In the top row, the plotted values are sampled at a selected longitude (ϕ=italic-$\phi=$180 degrees). In the bottom row, the same quantities but averaged over the longitude are shown.
@@
In Figure 6(f), the distributions of the averaged angular rotation rate are more similar to those inferred from observations: faster (prograde) equator rotation near the surface and slower (retrograde) high-latitude motions. The difference between the rotation rates at the solar equator and middle latitude region (±$\pm 30^{\circ}$ from the equator) is about 1.5 degrees per day or 40 degrees per one solar rotation, which is much larger than that of the actual Sun.
3.4 Divergence and vorticity of plasma flows
@@
The present simulation solves the fully compressible HD equations; hence, no assumptions are given to the divergence of the plasma flow, $\nabla\cdot{\bf{V}}$, or the mass flux, $\nabla\cdot(\varrho{\bf{V}})$. Figure 7 shows the latitude-longitude distribution of plasma flows near the top boundary surface. In panels (a) and (b), the divergence of the mass flux and the radial component of the vorticity of the plasma flow vector, $\nabla_{h}\times{\bf{V}}_{h}$, are shown, respectively. The convergence (negative divergence) of the mass flux is found at the downward flow (negative $V_{r}$) or the boundary of the large-scale granulation cells.
Figure 7: Latitude-longitude plot at $r=0.982\,R_{\sun}$ and $t=10^{7}$ s. The colors are truncated at the largest absolute value so that the red (for positive values) and blue (for negative values) are assigned evenly across the zero values.
@@
In Figure 7(b), the distribution of the radial vorticity exhibits several interesting features. First, the north-south symmetry is noticeable. Second, in the northern hemisphere, the positive (×$(\nabla\times{\bf{V}})_{r}$ values are concentrated along the boundaries of the convection cells, while the negative values are found rather evenly within the cells. Because the solar-like differential rotation yields the positive (×$(\nabla\times{\bf{V}})_{r}$ in the northern hemisphere, the obtained solar-like differential rotation (shown in panels (c) and (e) of Fig.6) are indeed contributed from the motions of the narrow downward-flow lanes. We think that the concentration of the positive radial vorticity is indeed formed in the divergent flows within each convection cell where the Coriolis force initiates the clockwise motions (viewed from outside) and forms lanes of the large positive vorticity (local counter-clockwise vortices).
Figure 8: The divergence of mass flux ($\nabla\cdot(\varrho{\bf{V}}$) and the normalized values in the meridional plane.
Figure 8(a) shows the meridional view of the divergence of the mass flux, $\nabla\cdot(\varrho{\bf{V}})$, averaged in the longitudinal direction. To analyze and quantify the divergence and its spatial distributions, we applied normalization to the divergences. In panel (b), the average of the absolute divergence, $|\nabla\cdot(\varrho{\bf{V}})|$, is shown.
@@
In Figure 8(c), the average of the absolute values is normalized with the average of the mass density, $|\nabla\cdot(\varrho{\bf{V}})|/\left<\varrho\right>$, so that the normalized value will be in the physical unit of ${\rm s}^{-1}$, and the inverse corresponds to the time scale for the divergence of the mass flux to substantially alter the local density. In this case, the larger values are found near the outer boundary surface, meaning that the relative variations of the plasma density are larger in the outer part of the convection zone than near the bottom of the convection zone, which is rather anticipated. The smallest value of the time scale (or the inverse of the largest value of the normalized divergence) is found near the top boundary surface and estimated at about $1.6\cdot 10^{5}s$ or about 1.8 $d$. The time scale for the deeper layers is estimated at an order of $10^{4}\,{\rm d}$. These estimations of time scale imply that the simulations with the divergence-free mass flux can properly handle the simulated system at least for this limit in a spherical convective shell, which does not include near-surface and surface layers. Of course, the size and position of the convective granulation changes quickly, so the practical limit must be much longer.
Figure 8(d) offers the distribution of $|\nabla\cdot(\varrho{\bf{V}})|/\left<\varrho V\right>$. This quantity has a physical unit of $\rm{cm}^{-1}$, and the inverse represents the spatial scale of the contribution of the divergence of the mass flux. The largest values are found near the top boundary, and the maxima is about $2.5\cdot 10^{-10}\,{\rm cm}^{-1}$, corresponding to the spatial scale of about $4\cdot 10^{9}{\rm cm}$, comparable to the thickness of the simulated convection zone. Near the bottom boundary surface, the spatial scale is estimated to be larger. The estimated spatial scales suggest that the discrepancy between the incompressible model and compressible one is overall not significant for the global scale of the solar convection simulations with the actual solar parameters.
@@
3.5 Spatial scale in the latitudinal and longitudinal directions
Panels (a) – (d) of Figure 9 show the spherical harmonic spectral powers of the three components of the plasma velocity ($V_{r}$, $V_{\theta}$, and $V_{\phi}$) and the kinetic energy density ($\varrho V^{2}/2$). Overall, the peaks are found around the harmonic degree ($l$) from 5 to 15 at any depths, and the peak values of these four parameters are larger (smaller) at the outer (inner) parts. The simulation grid size in the horizontal (perpendicular to the radial direction) is about 1 degree in the heliocentric angle, and a typical horizontal structure is captured by approximately 30 subgrid cells.
@@ 3.5 Spatial scale in the latitudinal and longitudinal directions
Figure 9: (a) – (d) Spherical harmonics spectral of three components of the plasma flow velocity ($V_{r}$, $V_{\theta}$, and $V_{\phi}$) and the kinetic energy density ($\varrho V^{2}/2$) at various depths. (e) The ratio of the power of $V_{\theta}$ (shown in panel (b)) to that of $V_{r}$ (shown in panel (a). (f) The ratio of the power of $V_{\phi}$ (shown in panel (c)) to that of $V_{r}$. In panels (e) and (f), the plot lines are truncated at $l=80$. The sampling depths are indicated with color thickness; the darker (lighter) blue for deeper (shallower) layers.
The powers for the longitudinal component $V_{\phi}$ and the kinetic energy ($\varrho V^{2}/2$) have peaks around $l\sim 5$, while the other two ($V_{r}$ and $V_{\theta}$) have peaks around $l\sim 10$. The powers of the plasma flow velocity become smaller as the sampling depth becomes deeper, while this tendency is not clear in the kinetic energy density.

### v3

@@ II.1 Model parameters, initial values, and simulation setup
Density at $r=R_{\rm bot}$, $\varrho_{\rm top}$ $g\cdot cm^{-3}]$ $2.1\cdot 10^{-1}$ $2.1\cdot 10^{-1}$
Density at $r=R_{\rm top}$, $\varrho_{\rm top}$ $g\cdot cm^{-3}]$ $2.1\cdot 10^{-4}$ $1.05\cdot 10^{-2}$
@@
kinematic viscosity $\nu$ $cm^{2}s^{-1}]$ $6.0\cdot 10^{12}$ $6.0\cdot 10^{13}$
entropy diffusion coef. $\kappa_{s}$ $cm^{2}s^{-1}]$ $6.0\cdot 10^{13}$ $6.0\cdot 10^{13}$
[…]
@@
Ekman num. $\nu/(\Omega_{o}D^{2})$ $10^{-14}$ $4.853\cdot 10^{-3}$ $2.447\cdot 10^{-3}$
Taylor num. $4\Omega_{o}^{2}D^{4}/\nu^{2}$ $10^{19\sim 27}$ $1.698\cdot 10^{5}$ $6.682\cdot 10^{5}$
Prandtl num. $\nu/\kappa_{s}$ $^{-6\sim-4} 0.10.1$ 11
Rossby num. $U^{\prime}_{\rm rms}/(2\Omega_{o}D)$ $10^{-1\sim 0}$ $1.239\cdot 10^{-1}$ $1.433\cdot 10^{-2}$
We set a lower value for the kinematic viscosity ($\nu=6\cdot 10^{12}$) than that in Chen et al. (2023) ($\nu=6\cdot 10^{13}$) to make the simulated system more closely match the actual Sun. Nonetheless, the scale parameters, such as the Rayleigh number and the Reynolds number, are different from those for the solar numbers by several to ten orders of magnitude. The chosen kinematic viscosity is the smallest value with which the simulation could run without numerical instability. Finer mesh assignments will allow us to simulate with a smaller value of the coefficient for the second-order derivatives; however, the use of such realistic viscosity or diffusivity would be left for future work. Given the current computational resources available, we set only the total luminosity and the rotation rate to those of the actual Sun.
The initial values are determined in the same manner as described in Chen et al. (2023). In brief, the initial velocity is set to zero, and the plasma density, gas pressure, and temperature are determined such that the polytropic relation, the heat flux conservation, and the balance between the radial pressure gradient and gravity will be satisfied. The calculation of the initial values requires the thermal heat conduction rate ($\kappa_{r}$) as a function of the radius and the density scale ($N_{p}$; the ratio of the mass density at the upper and lower boundary surfaces). We used the same values of the thermal heat conduction rate as in Chen et al. (2023) but a different value of the density scale (because of the difference in the radius of the top boundary surface, $R_{\rm top}$).
Figure 1: Radial profile of the simulated plasma quantities in the solar convective shell at $t=1\cdot 10^{7}s$: (a) mass density $\varrho$ and (c) temperature ($T$). In panels (b) and (d), the deviations of the simulated variable from the initial values, $\varrho_{0}$ and $T_{0}$, are given as $(\varrho/\varrho_{0}-1)$ and $(T/T_{0}-1)$, respectively. The deviations from the initial radial profile are less than $3\cdot 10^{-5}$ at $r<0.9R_{\sun}$ and about $2\cdot 10^{-3}$ near the top boundary. The contrast of the plasma density is tiny as $|\varrho/\left<\varrho\right>-1|<3\cdot 10^{-4}$, where $\left<\varrho\right$> is the averaged density over the spherical layer. Panels (e) and (f) show the relative deviation of the mass density and temperature from the averages, $(\varrho-\left<\varrho\right>)/\left<\varrho\right$> and $(T_{p}-\left<T_{p}\right>)/\left<T_{p}\right$>, respectively, with four lines representing the maximum, +1 standard deviation, -1 standard deviation, and the minimum values at each radius. In the bottom row, the radial profile of the square of buoyancy frequency ($N^{2}$) of the initial value and the buoyancy frequency averaged horizontally (latitudinally and longitudinally) are given in panels (e) and (f), respectively. In these two plots, cross (+) marks are placed where $N^{2}>0$.
@@
The initial state is nearly stable except in the near-surface regions at $r\sim R_{\rm top}$ where the plasma structure is unstable to perturbations or fluctuations. Figure 1(g) shows the radial profile of the square of the buoyancy frequency at the initial value,
@@
Figure 2 demonstrates the radial component of the plasma velocity ($V_{r}$) simulated at $t=1\cdot 10^{7}s$. The convective flows are seen in the simulated convection zone, as well as the convective cell structures on the top boundary surface. In Figure 1(h), the square of the buoyancy frequency $N^{2}$ is negative nearly all over the simulation volume at the selected moment.
Panels (a) and (c) of Figure 1, respectively, show the radial profile of the mass density and temperature averaged over each shell layer at $t=1\cdot 10^{7}$s. Panels (b) and (d), respectively, show the ratios of the simulated plasma density and temperature to those of the initial values. A notable difference is found in the density ratio near the top boundary surface, as shown in panel (b); still, the deviation is 0.25% in the largest. In panels (e) and (f) of Figure 1, the deviations of the plasma density and temperature from the average profile are small, too; the largest relative deviation of density and temperature, $(\varrho-\left<\varrho\right>)/\left<\varrho\right$> and $(T_{p}-\left<T_{p}\right>)/\left<T_{p}\right$>, are estimated at $\sim 5\cdot 10^{-4}$ and $\sim 1\cdot 10^{-4}$, respectively, except for the regions close to the 0.99-$R_{\sun}$ top boundary surface.
III Results
The kinetic energy density ($\varrho V^{2}/2$) is a good proxy to represent the entire simulated convective system. In Chen et al. (2023), the kinetic energy density of the simulated stellar convection zone exhibits exponential growth and a brief overshoot/oversaturation, then reaches a plateau profile that corresponds to the statistically stable saturated convection state.
@@ III Results
Figure 3 shows the temporal profile of the kinetic energy density. With the actual solar parameters, the temporal profile of the kinetic energy density is different from that in Chen et al. (2023). On the logarithmic scale plot, the exponential growth of the energy density is found as a (near) linear line, roughly at $10\leq t\leq 20$ days. However, the overshooting and the following plateau profile (like those in Figure 7 of Chen et al., 2023) are not clearly seen. The current simulation run has been performed only for $t\leq 1.5\cdot 10^{7}~{\rm s}\sim 170~{\rm d}$, but it is unclear whether or not the kinetic density will become more constant in an extended simulation run. Because our visual inspections did not find distinct differences in the figures for $t=1\cdot 10^{7}s$ and $t=1.5\cdot 10^{7}s$, we chose the simulated data at $t=10^{7}\,{\rm s}\sim 115\,{\rm d}$ as a moderately mature state of the convection properties.
III.1 Large-scale Convection Cells
@@
Figure 4: The plasma quantities of the simulated solar convection zone at $t=10^{7}$ s in the Mollweide projection mapping. (a) The plasma temperature subtracted by the average ($7.4132\cdot 10^{4}$ K). (b), (c), and (e) The radial, latitudinal, and longitudinal components of the plasma flow ($V_{r}$, $V_{\theta}$, and $V_{\phi}$), at $r=0.982\,R_{\sun}$. Panels (b), (d) and (f) in the right column compare the radial component at three depths, near the top boundary, at the middle depth of the simulated convection zone (at $r=0.846\,R_{\sun}$), and near the bottom boundary surface (at $r=0.709\,R_{\sun}$). The colors are truncated at the largest absolute value so that the red (for positive values) and blue (for negative values) are assigned evenly across the zero values.
@@
As shown in Figure 1(b,d), the deviations of the simulated plasma variables from the initial values are relatively small in the radial direction. The temperature, averaged over the selected sphere, was $7.4132\cdot 10^{4}\,{\rm K}$, approximately 8 degrees higher than the initial value, and the largest deviation from the averaged state is about 6.7 degrees. The lower temperature structures are found at the boundaries of the convection cells that can be identified as negative velocity $V_{r}$ in Figure 4(b). The temperature and plasma motions near the surface are found to be rather uneven within convection cell structures. For example, the region of relatively high temperature (panel a) and large prograde plasma motion ($V_{\phi}$ in panel e) is found in the western (right) part of the convection cells near the equator (panel b). The latitudinal motion ($V_{\theta}$ in panel c) exhibits a weak but similar tendency to $V_{\phi}$; relatively large poleward flows are found near the boundary of each convective cell close to the poles in each hemisphere. In the right column in Figure 4(b,d,f), the distribution of $V_{r}$ at different depths is presented. The size and shape of the convection cells differ substantially at different depths, while several strong downward flows are found at the same locations over a wide range of depths.
@@
Figure 5: (a) The radial component of the curl of plasma velocity (or the curl of the horizontal flows). (b) — (d) The $x$, $y$, and $z$ components of the simulated plasma flows around the north pole.
III.2 Polar regions
@@
Figure 5 offers the north pole view of the simulated plasma flows near the top boundary surface (at $r=0.982\,R_{\sun}$). Because the CHORUS++ code uses an unstructured grid system, issues due to the geometric singularity, like those in the case of the spherical coordinate formulation, are not present, and the plasma flows crossing the solar rotation axis (or $z$-axis) are treated naturally. The snapshot shown (at $t=1\cdot 10^{7}$ s) does not show clear polar vortex cell structures (e.g., Nagashima et al., 2011; Dikpati et al., 2024). By tuning our future simulation settings, we plan to investigate in more detail the formation of plasma-flow structures in the high-latitude and polar regions, which are believed to play a major role in solar-cycle activities.
Figure 6: The simulated plasma flows on the meridional plane at $t=10^{7}s$. From left, the radial component of the plasma flow ($V_{r}$), the latitudinal component ($V_{\theta}$), and the rotation rate relative to the reference frame ($\omega:=V_{\phi}/(r\sin\theta)$) are shown. In the top row (a) – (c), the values are sampled at a selected longitude ($\phi=$180 degrees). In the bottom row (d) – (f), the average over the longitude are shown.
III.3 Meridional Circulation and Rotation
@@
In Figure 6(f), the distributions of the averaged angular rotation rate are more similar to those inferred from observations: faster (prograde) equator rotation near the surface and slower (retrograde) high-latitude motions. The difference between the rotation rates at the solar equator and middle latitude region ($\pm 30^{\circ}$ from the equator) is about 1.5 degrees per day or 40 degrees per one solar rotation, which is much larger than that of the actual Sun.
Figure 7: Latitude-longitude plots of (a) the divergence of mass flux ($\nabla\cdot(\varrho\vec{V})$), (b) the radial component of the curl of plasma velocity ($\hat{r}\cdot(\nabla\times\vec{V})$), (c) the temperature subtracted by the average ($7.4132\cdot 10^{4}$ K), and (d) – (f) the radial, latitudinal, and longitudinal components of the plasma velocity ($V_{r}$$V_{\theta}$$V_{\phi}$) at $r=0.982\,R_{\sun}$ and $t=10^{7}$ s. The distributions of the spatial derivatives shown in panels (a) and (b) appear correlated with that of $V_{r}$ in panel (d). The colors are truncated at the largest absolute value so that the red (for positive values) and blue (for negative values) are assigned evenly across the zero values.
III.4 Divergence and vorticity of plasma flows
@@
The present simulation solves the fully compressible HD equations; hence, no assumptions are given to the divergence of the plasma flow, $\nabla\cdot\vec{V}$, or the mass flux, $\nabla\cdot(\varrho\vec{V})$. Figure 7 shows the latitude-longitude distribution of plasma flows near the top boundary surface. In panels (a) and (b), the divergence of the mass flux and the radial component of the vorticity of the plasma flow vector, $\nabla_{h}\times\vec{V}_{h}$, are shown, respectively. The convergence (negative divergence) of the mass flux is found at the region of the downward flow (negative $V_{r}$) and lower temperature, or the boundary of the large-scale convection cells.
@@
In Figure 7(b), the distribution of the radial component of the vorticity $(\nabla\times\vec{V})_{r}$ exhibits interesting features. In the northern hemisphere, for example, the positive (counter-clockwise vortices viewed from outside) values are concentrated along the boundaries of the convection cells (where $V_{r}<0$), while rather negative values are found evenly within the cells over the hemisphere. In the southern hemisphere, the same tendency, except for the signs of the vector quantities, is found. In the northern (southern) hemisphere of the examined state, the Coriolis force initiates the clockwise (counterclockwise) motions of the divergent flows of the convection cells, and the positive (negative) vorticity is formed where two negative(positive) vorticity regions contact each other. In panels (c) and (f) of Fig.6, we obtained a solar-like differential rotation pattern near the solar surface at the examined system state. This is because the positive vorticity is dominant over the negative one. It is necessary to conduct the simulation for a longer time to assess how long such unbalanced vorticity can persist and whether or not the vorticity imbalance can be found in a mature stage of global convection.
[…]
@@
To assess the differences between the presented model and anelastic ones, we calculate the divergence of $(\left<\varrho\right>\vec{V})$, where $\left<\varrho\right$> is the density averaged over each sphere of the same heliocentric distance. Figure 8(a) shows the meridional view of the divergence of the mass flux averaged over the longitude. In panel (b), the average of the absolute divergence, $|\nabla\cdot(\left<\varrho\right>\vec{V})|$, is shown. In Figure 8(c), the average of the absolute values is normalized with the average of the mass density, $|\nabla\cdot(\left<\varrho\right>\vec{V})|/\left<\varrho\right$>, so that the normalized value will be in the physical unit of ${\rm s}^{-1}$, and the inverse corresponds to the time scale for the divergence of the mass flux to substantially alter the local density. In this case, the larger values are found near the outer boundary surface, meaning that the relative variations of the plasma density are greater in the outer part of the convection zone than near the bottom of the convection zone, which is rather anticipated. The smallest value of the time scale (or the inverse of the largest value of the normalized divergence) is found near the top boundary surface and estimated at about $1.6\cdot 10^{5}$ seconds or about 1.8 days. The time scale for the deeper layers is estimated to be of an order of $10^{4}\,{\rm days}$. These time-scale estimates imply that simulations with a divergence-free mass flux can properly handle the simulated system at least for this limit in a spherical convective shell, which does not include near-surface and surface layers. Of course, the size and position of the convection cell change quickly, so the practical limit must be much longer.
[…]
@@
Figure 9: (a) – (d) Spherical harmonics spectral of three components of the plasma flow velocity ($V_{r}$, $V_{\theta}$, and $V_{\phi}$) and the kinetic energy density ($\varrho V^{2}/2$) at various depths. (e) The ratio of the power of $V_{\theta}$ (shown in panel (b)) to that of $V_{r}$ (shown in panel (a)). (f) The ratio of the power of $V_{\phi}$ (shown in panel (c)) to that of $V_{r}$. In panels (e) and (f), the plot lines are truncated at $l=80$. The sampling depths are indicated with color brightness; the darker (lighter) blue for deeper (shallower) layers.
III.5 Spatial scale in the latitudinal and longitudinal directions
Panels (a) – (d) of Figure 9 show the spherical harmonic spectral powers of the three components of the plasma velocity ($V_{r}$, $V_{\theta}$, and $V_{\phi}$) and the kinetic energy density ($\varrho V^{2}/2$). Overall, the peaks are found around the harmonic degree ($l$) from 5 to 15 at any depths, and the peak values of these four parameters are larger (smaller) at the outer (inner) parts. The simulation grid size in the horizontal (perpendicular to the radial direction) is about 1 degree in the heliocentric angle, and a typical horizontal structure is captured by approximately 30 subgrid cells. The powers for the longitudinal component $V_{\phi}$ and the kinetic energy ($\varrho V^{2}/2$) have peaks around $l\sim 5$, while the other two ($V_{r}$ and $V_{\theta}$) have peaks around $l\sim 10$. The powers of the plasma flow velocity become smaller as the sampling depth becomes deeper, while this tendency is not clear in the kinetic energy density.